# Swarm-intelligent search for gravitational waves from eccentric binary mergers

### v3

@@ II.4 Significance of candidates
When the iterations finish, the algorithm returns the final g-best template point with the associated triggers from individual detectors. The g-best template is already consistent in source parameters over all the detectors. Now these triggers are tested for any zero-lag time coincidence to create foreground events. Several such foreground candidates, which pass both the parameter and the time coincidence tests, may still not be of astrophysical origin. To this end, one can assign a statistical significance to a foreground candidate and/or a probability of being of astrophysical origin, (Farr et al., 2015; Capano et al., 2017). Here we assign a false alarm rate, FAR to the foreground events using the widely used method of time-shifting of triggers obtained from processing the data immediately available before the candidate. If triggers obtained from a pair of detectors are artificially shifted with respect to each other by a time more than travel time of light between the two detectors, then the resulting coincidences cannot be of astrophysical origin. Such artificial background events are also ranked by their artificial network statistic value and a FAR estimate is assigned to a foreground candidate as
@@
Here is the number of background events with a network statistic equal to or greater than the foreground event. is computed as the artificial time over which the background coincident events are manufactured. In the simple case of a two detector network, it can be approximated by , where is the coincident live time of the two detectors and is a non-astrophysical time-slide duration. To compute a multi-detector background, the triggers from the most sensitive detector may be slided with remaining detectors fixed (Davies et al., 2020). Even considering triggers from Virgo, a time-slide duration of 50 ms can be safely chosen and as many artificial events are created as possible. When additional detectors are ready with similar sensitivities, a general approach with each detector sliding may be appropriate. Thus, to attempt a significant confidence of FAR less than 1 per 100 years to a given foreground event, a buffer of a sufficient duration of recent triggers must be kept which turns out to be a minimum of 3.5 hrs in our case. Normally, a large window for buffering triggers is more exposed to less recent and abrupt changes in the detector noise. Online LVK searches, however, require a higher FAR threshold for considering candidates for follow-up processes. In this work, we have mostly used only 4096 seconds to reach FARs upto 1 per 10 years.
@@ III Eccentric Search
In the previous section, we described a general framework to search for CBC signals using the PSO algorithm. Now, we demonstrate the algorithm specifically designed for spin-aligned eccentric sources. In the absence of the occasional and loud disturbances, or glitches, the real instrument data can be considered as nearly Gaussian. Thus in this section, we first demonstrate the recovery of simulated signals with our algorithm in coloured Gaussian noise. The signals are projected from random sky locations onto a network of simulated detectors and then injected into the noise segments. Here we at most use the 3-detector combination, from H1, L1, and V1, and use the publicly available O4 representative PSDs to colour the noise segments (https://dcc.ligo.org/LIGO T2000012/public, ). We then discuss the response of the search to the real data from the instruments when no likely signals are present. With each of these, we also include the corresponding observations for the aligned-spin searches. To reduce the cost of matched-filtering, data are downsampled to 1024 Hz while remaining sensitive to almost the full inspiral part of the signal as shown in the Fig. 5. Note that the TaylorF2Ecc is an inspiral-only model and does not anyway contain much power in the high frequencies. The loss in the SNR from such a high frequency cutoff would be just around 0.2 for a binary and even smaller for heavier binaries. We currently set the low frequency cutoff at 20 Hz. Since, the effects of eccentricity are more in the lower frequencies, it is reasonable to expect improvements with a lower low frequency cutoff.
@@ III.1 Eccentric BNSs
In this section, we generate mock signals from a simulated population of about 500 eccentric BNS sources using the TEOBResumS-DALI model and inject them into the Gaussian noise simulated for the two LIGO detectors. Agnostic about the true astrophysical distribution, the merger component masses are uniformly drawn from , with moderately large values of the aligned-spin components in and small eccentricities from . We use TaylorF2Ecc to recover the signals using the PSO algorithm. The TaylorF2Ecc waveform model basically draws the parameter space coverage from the widely used TaylorF2 model to search for BNS systems. For long duration signals, we expect to see an insignificant SNR loss due to the missing merger-ringdown parts in the template waveforms as shown in (Fig. 6). We test the accuracy of the PSO algorithm to mainly recover the coincident SNR and the injected binary parameters. Note that our primary goal here is to demonstrate the effectiveness of the PSO algorithm, but it also tests the agreement of the waveform parameters between the two models. The choice of the injection distribution may not be astrophysically well informed, however, the luminosity distances to the simulated sources are chosen from Mpc distributed uniformly in volume, giving a realistic range of network SNRs for O4 and beyond. In Fig. 8, the accuracy in recovering the injected parameters is plotted for the component masses, the effective spin parameters and the eccentricity. We find that the PSO algorithm is readily sensitive to these small eccentrities with the total 75k matched-filtering operations per detector. Note that, even a small eccentricity can modulate the signal over a large number of cycles, attributing to the improvement in the recovered coincident SNR as shown in the Fig. 6. It also may be concluded that the waveform parameter values are largely precise for the two models from different formalisms. Note that whatever deviation observed here also includes the fact that the templates do not span the merger-ringdown phases of the signals.
@@ III.2 Eccentric BBHs
Here we use the TEOBResumS-DALI waveform model for both the signal injection and the template generation. This waveform model contains the inspiral-merger-ringdown phases required for the SNR build up, especially for the high mass binaries. As before, Fig. 9 summarizes the recovered parameters for the H1 and L1 coincidences out of 1000 injections in the mass range , aligned-spins within a range of and eccentricities up to . The luminosity distances of the sources are chosen uniformly in volume between 500-3000 Mpc. The search was performed using about 30k matched-filtering operations per detector.
@@ III.3 Background event rates
In the previous sections, we have shown that the eccentric searches can outperform the aligned-spin searches in the presence of Gaussian noise. However, the data from a real detector can contain nearby noise glitches which can overwhelm the searches. To investigate this, we start with an hypothesis that there is no strong astrophysical signal and data segments contain only detector noise and glitches. We perform the searches independently using same configuration as described in section III on several data chunks from the third observing run (O3a) of the Advanced LIGOs not containing any obvious CBC candidates (Abbott et al., 2023, 2021e). For simplicity we use data from only two detectors LHO and LLO. Note that the triggers crossing an SNR threshold undergo the same signal consitency tests and the network reweighted-SNR is used as the ranking statistic. For comparison, we also conduct the usual aligned-spin search on the same data segments with identical set-up. The results are summarized in the Fig. 11.
@@
While searching for the BNS systems, we find that both the searches generate a similar rate of background events. The power -test is effective in this low mass region with long inspirals and vetoes out most of the glitches. On the other hand, the eccentric BBH search uses a time domain model for the templates, where the filter artefacts can lead to higher rate of background events if not conditioned properly. We constrain the waveform with the tapering scheme for time-domain templates illustrated in (McKechan et al., 2010). In addition, it is known that some of the glitches, such as the blips, can successfully dodge the power -test for the higher-mass templates. In the BBH searches, the component masses can go beyond and eccentricity can further shorten the template durations. Sine-gaussians can approximately model some of these glitches and separate them from GW signals by taking into account of the power at frequencies higher than those in the templates (Nitz, 2018). To benefit and extend the eccentric BBH search for high mass systems, this can be incorporated in the future when we expand the search space coverage.

### v5

@@ II.4 Significance of candidates
When the iterations finish, the algorithm returns the detection template for a given stretch of data with the associated triggers from the individual detectors. The detection template point is already consistent in the source parameters across all the detectors. Now the secondary triggers are tested for any zero-lag time coincidence to create foreground events. A time coincidence test requires triggers from different detectors to occur within a time window defined by the light’s travel time between the detectors plus 5 ms accounting for any possible timing measurement errors. Several such foreground candidates, which pass both the parameter and the time coincidence tests, may still not be of astrophysical origin. To assess the possibility of a noise origin, one assigns a statistical significance to a foreground candidate and/or a probability of being of astrophysical origin, (Farr et al., 2015; Capano et al., 2017). Here we assign a false alarm rate (FAR) to the foreground events using the method of time-shifting of triggers obtained from processing the data immediately available near (or before) the candidate (Wąs et al., 2009). The algorithm maintains a sufficient buffer of the cumulative secondary triggers resulting from the detection templates of the neighbouring (or past) data stretches. Note that the different stretches can have their optimized g-best template points located throughout the parameter space. Thus, though the secondary triggers for a given data stretch have consistency over the source parameters, the cumulative of such triggers over multiple stretches do not enforce the source parameter consistency 222An appropriate strategy to utilize even the primary triggers for the background estimates in the PSO search can be explored in a future study. However, the cumulative triggers from the entire space are generally more prone to picking up abrupt noise and can even degrade the background statistic.. If the triggers obtained from a pair of detectors are artificially shifted with respect to each other by a time more than travel time of light between the two detectors, then the resulting coincidences cannot be of astrophysical origin. Such artificial background events are also ranked by their artificial network statistic value and a FAR estimate is assigned to a foreground candidate as
@@
Here is the number of background events with a network statistic equal to or greater than the foreground event. is computed as the artificial time over which the background coincident events are manufactured. In the simple case of a two detector network, it can be approximated by , where is the coincident live time of the two detectors and is a non-astrophysical time-slide duration. To compute a multi-detector background, the triggers from the most sensitive detector may be slided with remaining detectors fixed (Davies et al., 2020). Even considering triggers from Virgo, a time-slide duration of 50 ms can be safely chosen and as many artificial events are created as possible. When additional detectors are ready with similar sensitivities, a general approach with each detector sliding may be appropriate. So, to attempt a significant confidence of FAR less than 1 per 100 years to a given foreground event, a buffer of a sufficient duration of recent triggers must be kept which turns out to be a minimum of 3.5 hrs in our case. Normally, a large window for buffering triggers is more exposed to less recent and abrupt changes in the detector noise. Online LVK searches, however, require a higher FAR threshold for considering candidates for follow-up processes. In this work, we have mostly used only 4096 seconds to reach FARs upto 1 per 10 years.
@@ III Eccentric Search
In the previous section, we described a general framework to search for CBC signals using the PSO algorithm. Now, we demonstrate the algorithm specifically designed for spin-aligned eccentric sources. Here the search dimensions are given by
We use a uniform initial distribution of particles in the added eccentric dimension, . In the absence of the occasional and loud disturbances, or glitches, the real instrument data can be considered as nearly Gaussian. Thus in this section, we first demonstrate the recovery of simulated signals with our algorithm in coloured Gaussian noise. The signals are projected from random sky locations onto a network of simulated detectors and then injected into the noise segments. Here we at most use the 3-detector combination, from H1, L1, and V1, and use the publicly available O4 representative PSDs to colour the noise segments (https://dcc.ligo.org/LIGO T2000012/public, ). We then discuss the response of the search to the real data from the instruments when no likely signals are present. For each set of results obtained with the aligned-spin eccentric search using PSO, we also show results from a corresponding aligned-spin non-eccentric search with PSO.
To reduce the cost of matched-filtering, data are downsampled to 1024 Hz while remaining sensitive to almost the full inspiral part of the signal as shown in Fig. 5. Note that the TaylorF2Ecc is an inspiral-only model and does not anyway contain much power in the high frequencies. The loss in the recovered SNR from such a high frequency cutoff would be just around 1% in the case of a perfectly matching template for a binary and even smaller for heavier binaries. We currently set the low frequency cutoff at 20 Hz. Since, the effects of eccentricity are more in the lower frequencies, it is reasonable to expect improvements with a lower low frequency cutoff.
@@ III.1 Eccentric BNSs
In this section, we generate mock signals from a simulated population of about 500 eccentric BNS sources using the TEOBResumS-DALI model and inject them into the Gaussian noise simulated for the two LIGO detectors. Agnostic about the true astrophysical distribution, the merger component masses are uniformly drawn from , with moderately large values of the aligned-spin components in and small eccentricities from . We use TaylorF2Ecc to recover the signals using the PSO algorithm. The TaylorF2Ecc waveform model basically draws the parameter space coverage from the widely used TaylorF2 model to search for BNS systems. For long duration signals, we expect to see an insignificant SNR loss due to the missing merger-ringdown parts in the template waveforms as shown in (Fig. 6). We test the accuracy of the PSO algorithm to mainly recover the coincident SNR and the injected binary parameters. Note that our primary goal here is to demonstrate the effectiveness of the PSO algorithm, but it also tests the agreement of the waveform parameters between the two models. The choice of the injection distribution may not be astrophysically well informed, however, the luminosity distances to the simulated sources are chosen from Mpc distributed uniformly in volume, giving a realistic range of network SNRs for O4 and beyond. In Fig. 9, the accuracy in recovering the injected parameters is plotted for the component masses, the effective spin parameters and the eccentricity. We find that the PSO algorithm is readily sensitive to these small eccentrities with the total 75k matched-filtering operations per detector per injection. Note that, even a small eccentricity can modulate the signal over a large number of cycles, attributing to the improvement in the recovered coincident SNR as shown in the Fig. 6. It also may be concluded that the waveform parameter values are largely precise for the two models from different formalisms. Note that whatever deviation observed here also includes the fact that the templates do not span the merger-ringdown phases of the signals.
@@ III.2 Eccentric BBHs
Here we use the TEOBResumS-DALI waveform model for both the signal injection and the template generation. This waveform model contains the inspiral-merger-ringdown phases required for the SNR build up, especially for the high mass binaries. As before, Fig. 9 summarizes the recovered parameters for the HL coincidences out of 1000 injections in the mass range , aligned-spins within a range of and eccentricities up to . The luminosity distances of the sources are chosen uniformly in volume between 500-3000 Mpc. The search was performed using about 30k matched-filtering operations per detector per injection.
@@ III.3 Background event rates
Event | Triggered | GPS time | FAR | |||||||||
(seconds) | (year-1) | PSO | PyCBC | MBTA | GstLAL | |||||||
GW150914 | HL | 1126259462.4 | 46.1 | 26.3 | 0.002 | 0.073 | 0.38 | 23.7 | 23.7 | 23.6 | - | 24.4 |
GW151012 | HL | 1128678900.4 | 26.6 | 17.6 | 0.099 | 0.283 | 0.38 | 9.7 | 9.8 | 9.5 | - | 10 |
GW170104 | HL | 1167559936.5 | 40.4 | 20.7 | -0.034 | 0.035 | 0.38 | 13.1 | 13.2 | 13 | - | 13 |
GW170729 | HL | 1185389807.3 | 70.1 | 29.1 | 0.051 | 0.325 | 2.26 | 9.5 | 9.9 | 9.8 | 10.8 | |
GW170809 | HL | 1186302519.7 | 34.7 | 27.2 | -0.186 | 0.277 | 0.38 | 10.2 | 10.8 | 12.2 | - | 12.4 |
GW170814 | HL | 1186741861.5 | 37.5 | 26.9 | 0.179 | 0.2 | 0.38 | 12.7 | 15.0 | 16.3 | - | 15.9 |
GW170817@ | HL | 1187008882.4 | 1.66 | 1.15 | -0.001 | 0.01 | 0.38 | 32.4 | 32.5 | 30.9 | - | 33.0 |
GW170818 | HL | 1187058327.1 | 44.8 | 22.1 | -0.185 | 0.218 | 0.38 | 9.9 | 10.0 | - | - | 11.3 |
GW170823 | HL | 1187529256.5 | 51.2 | 43.8 | 0.264 | 0.056 | 0.38 | 11.3 | 11.6 | 11.1 | - | 11.5 |
GW190403_051519 | - | - | - | - | - | - | - | - | - | - | - | - |
GW190408_181802 | HL | 1238782700.3 | 41.9 | 17.2 | 0.002 | 0.104 | 0.38 | 14.1 | 14.1 | 13.1 | 14.4 | 14.7 |
GW190413_052954 | HL | 1239168612.5 | 54.9 | 43.6 | 0.126 | 0.071 | 0.38 | 8.7 | 8.7 | 8.5 | - | - |
GW190413_134308 | HL | 1239198206.7 | 82.6 | 37.6 | -0.297 | 0 | 0.38 | 9.2 | 9.5 | 9.3 | 10.3 | 10.1 |
GW190421_213856 | HL | 1239917954.2 | 63.5 | 40.8 | -0.148 | 0.108 | 0.38 | 10.2 | 10.2 | 10.1 | 9.7 | 10.5 |
GW190503_185404 | HL | 1240944862.3 | 49.4 | 34.6 | -0.097 | 0.204 | 0.38 | 12.2 | 12.2 | 12.2 | 12.8 | 12 |
GW190512_180714 | HL | 1241719652.4 | 27.5 | 15.9 | -0.068 | 0.12 | 0.38 | 11.4 | 11.6 | 12.4 | 11.7 | 12.2 |
GW190513_205428 | - | - | - | - | - | - | - | - | - | 11.6 | 13 | 12.3 |
GW190514_065416 | - | - | - | - | - | - | - | - | - | - | - | 8.4 |
GW190517_055101 | HL | 1242107479.8 | 54.4 | 24.9 | 0.365 | 0.306 | 0.38 | 9.5 | 9.5 | 10.4 | 10.8 | 11.3 |
GW190519_153544 | HL | 1242315362.4 | 92.6 | 85.3 | -0.499 | 0.191 | 2.63 | 11.3 | 11.3 | 13.2 | 13.7 | 12.4 |
GW190521_030229∗ | HL | 1242442967.4 | 143.5 | 133.8 | 0.223 | 0.246 | 0.009 | 13.5 | 13.7 | 13.7 | 13 | 13.3 |
GW190521_074359 | HL | 1242459857.4 | 51.8 | 43.2 | 0.162 | 0.005 | 0.38 | 24.2 | 24.4 | 24 | 22.2 | 24.4 |
GW190527_092055 | HL | 1242984073.8 | 60.7 | 25.1 | 0.088 | 0.322 | 0.38 | 9.0 | 9.0 | - | - | 8.7 |
GW190602_175927 | HL | 1243533585.1 | 93.1 | 72.0 | 0.036 | 0.195 | 0.38 | 12.2 | 12.2 | 11.9 | 12.6 | 12.3 |
GW190620_030421 | - | - | - | - | - | - | - | - | - | - | - | 10.9 |
GW190630_185205 | - | - | - | -. | - | - | - | - | - | - | 15.2 | - |
GW190701_203306 | HLV | 1246048404.6 | 69.9 | 56.3 | -0.235 | 0.436 | 2.63 | 12.3 | 12.3 | 11.9 | 11.3 | 11.7 |
GW190706_222641 | HL | 1246487219.3 | 90.3 | 48.7 | 0.148 | 0.497 | 0.38 | 13.0 | 13.0 | 11.7 | 11.9 | 12.5 |
GW190708_232457 | - | - | - | - | - | - | - | - | - | - | - | 13.1 |
GW190719_215514 | HL | 1247608532.9 | 53.9 | 30.6 | 0.184 | 0.164 | 0.38 | 8.5 | 8.5 | - | - | - |
GW190727_060333 | HL | 1248242631.9 | 53.3 | 48.5 | 0.151 | 0.007 | 0.38 | 11.7 | 11.7 | 11.4 | 12 | 12.1 |
GW190731_140936 | HL | 1248617394.6 | 66.1 | 16.7 | -0.356 | 0.09 | 0.38 | 8.6 | 8.6 | - | 9.1 | 8.5 |
GW190803_022701 | HL | 1248834439.8 | 58.6 | 36.7 | -0.119 | 0.292 | 5.64 | 8.2 | 8.5 | 8.7 | 9 | 9.1 |
GW190805_211137 | HL | 1249074715.3 | 88.5 | 35.2 | 0.326 | 0.335 | 0.38 | 7.9 | 7.9 | - | - | - |
GW190828_063405 | HL | 1251009263.7 | 39.6 | 36.1 | 0.169 | 0.164 | 0.38 | 15.7 | 15.7 | 13.9 | 15.2 | 16.3 |
GW190828_065509 | HL | 1251010527.8 | 29.7 | 13.9 | 0.075 | 0.158 | 0.75 | 11.0 | 11.0 | 10.5 | 11.1 | 10.8 |
GW190910_112807 | - | - | - | - | - | - | - | - | - | - | - | 13.4 |
GW190915_235702 | HL | 1252627040.6 | 42.2 | 28.5 | -0.116 | 0.175 | 0.38 | 12.8 | 12.8 | 13 | 12.7 | 13 |
GW190916_200658 | HL | 1252699636.8 | 48.9 | 35.9 | 0.05 | 0.259 | 381.12 | 6.7 | 6.8 | - | 8.2 | 8.2 |
GW190925_232845 | HV | 1253489343.1 | 28.2 | 14.8 | 0 | 0.127 | 0.38 | 9.9 | 9.9 | 9 | 9.4 | - |
GW190926_050336 | HL | 1253509434.0 | 87.5 | 16.2 | -0.208 | 0.338 | 1.13 | 8.1 | 8.1 | - | - | 9 |
GW190929_012149 | HL | 1253755327.4 | 88.3 | 46.9 | -0.062 | 0.14 | 0.75 | 9.6 | 9.6 | 9.4 | 10.3 | 10.1 |
GW191109_010717 | HL | 1257296855.2 | 80.4 | 59.6 | 0.176 | 0.333 | 0.38 | 16.0 | 16.0 | 13.2 | 15.2 | 15.8 |
GW191127_050227 | HL | 1258866165.5 | 46.7 | 26.5 | -0.214 | 0 | 1.50 | 8.7 | 8.7 | 9.5 | 9.8 | 10.3 |
GW191204_110529 | HL | 1259492747.5 | 35.5 | 27.6 | 0.131 | 0.023 | 0.38 | 9.6 | 9.6 | 8.9 | 8.1 | 9 |
GW191215_223052 | HL | 1260484270.3 | 39.5 | 21.1 | 0.048 | 0.085 | 0.38 | 10.6 | 10.6 | 10.3 | 10.8 | 10.9 |
GW191222_033537 | HL | 1261020955.1 | 68.5 | 51.6 | 0.018 | 0.092 | 0.38 | 11.4 | 11.8 | 11.5 | 10.8 | 12 |
GW191230_180458 | HL | 1261764316.4 | 90.4 | 37.9 | 0.075 | 0.109 | 0.38 | 9.8 | 9.8 | 9.6 | 9.8 | 10.3 |
GW200112_155838 | - | - | - | - | - | - | - | - | - | - | - | 17.6 |
GW200128_022011 | HL | 1264213229.9 | 70.6 | 52.2 | 0.25 | 0.078 | 0.38 | 10.1 | 10.1 | 9.8 | 9.4 | 10.1 |
GW200129_065458 | HLV | 1264316116.4 | 40.4 | 31.4 | 0.156 | 0.201 | 0.38 | 20.8 | 25.5 | 16.3 | - | 26.5 |
GW200208_130117 | HLV | 1265202095.9 | 51.7 | 33.5 | -0.13 | 0.036 | 0.38 | 10.8 | 10.8 | 9.6 | 10.4 | 10.7 |
GW200208_222617 | - | - | - | - | - | - | - | - | - | - | 8.9 | 8.2 |
GW200209_085452 | HL | 1265273710.1 | 67.8 | 37.1 | -0.137 | 0.149 | 0.38 | 9.0 | 9.0 | 9.2 | 9.7 | 10 |
GW200216_220804 | HL | 1265926102.8 | 79.5 | 32.7 | 0.107 | 0.12 | 0.38 | 8.8 | 8.8 | 9 | 8.8 | 9.4 |
GW200219_094415 | HL | 1266140673.1 | 63.7 | 39.0 | -0.042 | 0 | 1.13 | 10.0 | 10.0 | 9.9 | 10.6 | 10.7 |
GW200220_061928 | - | - | - | - | - | - | - | - | - | - | - | - |
GW200220_124850 | - | - | - | - | - | - | - | - | - | - | 8.2 | 8.2 |
GW200224_222234 | HLV | 1266618172.4 | 49.6 | 36.8 | 0.105 | 0.169 | 0.38 | 14.8 | 18.1 | 19.2 | 19 | 18.9 |
GW200225_060421 | HL | 1266645879.3 | 23.0 | 18.7 | -0.037 | 0.088 | 0.38 | 11.1 | 12.3 | 12.3 | 12.5 | 12.9 |
GW200302_015811 | - | - | - | - | - | - | - | - | - | - | - | 10.6 |
GW200306_093714 | - | - | - | - | - | - | - | - | - | 7.8 | 8.5 | - |
GW200308_173609 | - | - | - | - | - | - | - | - | - | 7.9 | 8.3 | 8.1 |
GW200311_115853 | HLV | 1267963151.3 | 40.7 | 32.2 | -0.019 | 0.046 | 0.38 | 15.8 | 17.1 | 17 | 16.5 | 17.7 |
GW200322_091133 | - | - | - | - | - | - | - | - | - | 8 | 9 | - |
@@
In the previous sections, we have shown that the eccentric searches can outperform the aligned-spin searches in the presence of Gaussian noise. However, the data from a real detector can contain nearby noise glitches which can overwhelm the searches. To investigate this, we start with an hypothesis that there is no strong astrophysical signal and data segments contain only detector noise and glitches. We perform the searches independently using the same configuration as described in section III on several data blocks from the third observing run (O3a) of the Advanced LIGOs not containing any obvious CBC candidates (Abbott et al., 2023, 2021e). For simplicity we use data from only two detectors H1 and L1. Each 4096 seconds data block is divided into segments of 256 seconds with a 50% overlap, equally shared with its neighbouring segments and the analysis procedure described in Sec. II.3 is followed. Thus each PSO instance effectively searches 128 seconds of data from the available detectors. Note that the triggers crossing an SNR threshold undergo the same signal consitency tests and the network reweighted SNR is used as the ranking statistic. For comparison, we also conduct the aligned-spin non-eccentric search using PSO on the same data segments with identical set-up. The results are summarized in Fig. 11.
@@
While searching for the BNS systems, we find that both the searches generate a similar rate of background events. The power -test is effective in this low mass region with long inspirals and vetoes out most of the glitches. On the other hand, the eccentric BBH search uses a time domain model for the templates, where the filter artefacts can lead to higher rate of background events if not conditioned properly. We constrain the waveform with the tapering scheme for time-domain templates illustrated in (McKechan et al., 2010). In addition, it is known that some of the glitches, such as the blips, can successfully dodge the power -test for the higher-mass templates. In the BBH searches, the component masses can go beyond 100 and eccentricity can further shorten the template durations. Sine-gaussians can approximately model some of these glitches and separate them from GW signals by taking into account of the power at frequencies higher than those in the templates (Nitz, 2018). To benefit and extend the eccentric BBH search for high mass systems, this can be incorporated in the future when we expand the search space coverage.
However, to assess the accuracy of the estimation of the false alarm rates described earlier in subsection II.4, we use the triggers obtained from analyzing the above datasets. The cumulative number of all the foreground candidates are plotted against a given inverse false alarm rate (IFAR) threshold in Fig. 12. Here the FARs are assigned using the cumulative triggers from both BNS and BBH searches, inclusive of those forming the foreground candidates. The expected number of noise events for a given FAR threshold over a coincident observation time is calculated using Eq. 6. The figure demonstrates a general self consistency of the FARs assigned to any of the foreground candidates and indicates that the candidates are consistent with noise origin in this case.